OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y less than or greater to -x - 4 (1, 10) (-1, 10) (10, 1) (1, -10)

OpenStudy (bahrom7893):

Less than or greater to?

OpenStudy (anonymous):

<=

OpenStudy (phi):

that is less than or equal try each of the choices and see which one works

OpenStudy (anonymous):

\[y \le x-5 y \le -x-4\]

OpenStudy (anonymous):

but sepperate the 5 and the y

OpenStudy (phi):

yes, two separate relations y ≤ x - 5 and y ≤ -x -4

OpenStudy (phi):

I would test the pair with the smallest y as a first guess

OpenStudy (anonymous):

um okay I'll try...

OpenStudy (anonymous):

Can you explain how I fit the coordinates into the equation...?

OpenStudy (phi):

the points are (x,y) pairs to test if a point "works" replace x and y in the relation for example (1,10) has x=1 and y=10 plop those into the first relation y ≤ x - 5 10 ≤ 1 - 5 do the arithmetic 10≤ -4 no 10 is not less than or equal to -4 so the first choice is out. if it did work, we would test that the point also worked for the second relation

OpenStudy (anonymous):

Ohhh okay. Let me do the math real quick :)

OpenStudy (anonymous):

Is it B?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

Wait i think I did it wrong

OpenStudy (phi):

how did you do it ?

OpenStudy (anonymous):

Actually I think it's D because -10<= 1-5 is true and so it -10<=-1-5 because 4 is greater than -10 and so is -6

OpenStudy (phi):

yes, it is D. but you mean -10 ≤ 1-5 or -10 ≤ - 4 (not +4) I assume that was a typo

OpenStudy (anonymous):

Yes it was :)